# Groups with presentations in EDT0L

### 02 Feb 2024 v1

@@ 2.2.3 Tietze Transformations
Here, we give, thanks to Tietze transformations, natural conditions on a class of languages LAN so that the set of marked groups $\pi_{M}(\text{LAN})$ be saturated for the abstract group isomorphism relation. The conditions are given in Lemma 28.
@@
If LAN is a class of languages stable under taking homomorphic images and taking union with finite sets, then the set of marked groups that have a presentation in LAN is closed under group isomorphism: if a group has a presentation in LAN with respect to a marking, it has one with respect to any marking.
@@
Suppose that LAN is a class of uniformly recursively enumerable languages. Suppose furthermore that LAN is effectively stable under taking homomorphic images and taking union with finite sets. Then there is a procedure that takes as input a LAN-presentation of a marked groups $(G,S)$, and outputs a sequence of LAN-presentations of $G$ that contains at least one LAN-presentation for each marking of $G$.
@@
the hypotheses on the class LAN of languages immediately imply that this presentation is again a LAN-presentation. ∎
@@ 2.3 Computing quotients and dismissing infinite presentations
Let LAN be a class of languages, and $\mathcal{C}$ a set of finitely presented groups. In this section, we give conditions under which the following two problems are equivalent:
@@
Let $\mathcal{C}$ be a recursively enumerable class of finitely presented groups. Let LAN be a class of uniformly recursively enumerable languages. If the marked quotient problem in $\mathcal{C}$ is uniformly semi-decidable for groups given by presentations in LAN, then there is an algorithm which, given as input a presentation in LAN for a marked group in $\mathcal{C}$, produces a finite presentation for this group.
@@
An open problem is to find a class LAN of formal languages for which the above lemma can be applied with $\mathcal{C}$ being the set of all finitely presented groups. It is an open problem to decide whether groups given by EDT0L presentations have uniformly semi-decidable marked quotient problem for finitely presented groups.
@@
Let $\mathcal{C}$ be a class of finitely presented groups. Suppose that the isomorphism problem is solvable in $\mathcal{C}$ for groups given by finite presentations, and that $\mathcal{C}$ is recursively enumerable via finite presentations. Let LAN be class of languages. If the marked quotient problem in $\mathcal{C}$ is uniformly semi-decidable for groups given by presentations in LAN, then the isomorphism problem is solvable for groups in $\mathcal{C}$ given by presentations in LAN.
@@
Let $\mathcal{C}$ be a class of finitely presented groups stable under taking quotients. Let LAN be a class of languages for which the operation “taking union with finite sets” is computable. Then, if the marked isomorphism problem is solvable for groups in $\mathcal{C}$ with presentations in LAN, the $\mathcal{C}$-marked quotient problem is uniformly semi-decidable for groups given by LAN presentations.
@@ 3.1 Identification result
Proof of Theorem 37.
From the definitions it is obvious that L-presentations produce their relators exactly by the DTF0L+FIN mechanism, with an additional constraint: the iterated morphisms should be group endomorphisms, and not simply monoid endomorphisms. Thus having an L-presentation implies having a DTF0L+FIN presentation. And of course, since DTF0L+FIN is a proper subclass of EDT0L, having a presentation in DTF0L+FIN implies having a presentation in EDT0L.
@@ 3.2 Application to the marked isomorphism problem
We use the fact that equality of EDT0L languages is not semi-decidable, see [Roz72]. (Note that the theorem stated in [Roz72] is “equality is undecidable for EDT0L languages”. But for every class of uniformly recursive languages the equality problem is co-semi-decidable, since to prove that two languages are different it suffices to produce an element which belongs to one but not to the other. Hence the statement that equality is not decidable shows actually that equality is not semi-decidable for EDT0L languages.)
@@
One could expect much stronger results than the ones above. It would be very interesting to find examples of a class of languages LAN where the equality and inclusion problems are semi-decidable, while the isomorphism and quotient problem are not semi-decidable for groups with presentations in LAN.
@@
4 Unified approach to finite, nilpotent, metabelian, free and hyperbolic quotients
In this section we give a unified approach to computing quotients of EDT0L presented groups, thereby unifying and generalizing [BEH08] and [Har11]. (Note also that [BEH08] and [Har11] deal with L-presentations, and while this does not make a difference in terms of groups, it possibly does in terms of marked groups, and the following results deal with marked groups.) The following theorems were known:
@@ 4 Unified approach to finite, nilpotent, metabelian, free and hyperbolic quotients
Theorem 47 (Metabelian quotients).
There is an algorithm that, given an EDT0L presentation for a marked group $(G,S)$, computes a finite metabelian presentation for the group $(G/[[G,G],[G,G]],S)$.
@@
Theorem 49 (Hyperbolic quotients).
There is an algorithm that, on input an EDT0L presentation for a marked group $(G,S)$, a finite presentation for a marked hyperbolic group $(H,S^{\prime})$, and a function $f:S\rightarrow S^{\prime}$, decides whether or not $f$ can be extended to a group homomorphism.
The theorems above in each case rely on two properties (we separate the case of finite quotients from that of hyperbolic quotients to highlight the fact that its proof can rely on simpler arguments):
  • For nilpotent and finite groups, on noetherianity and uniform solvability of the word problem.
  • For metabelian groups, noetherianity is replaced by the Max-n property, which is noetherianity for normal subgroups. We also use uniform solvability of the word problem.
  • For hyperbolic groups, we rely on equational noetherianity instead of actual noetherianity, and on solvability of the universal theory of each hyperbolic groups ([DG10, Corollary 9.1]) instead of the word problem.
@@ 4.1 Subgroup functors of fully characteristic subgroups
A radical is a subgroup functor $\kappa$ such that:
@@
We will see several examples of subgroup functors that satisfy the desired conditions.
Example 53.
A general construction of functors that satisfy our conditions is as follows.
Consider a set $\mathcal{K}$ of groups. Then we can associate to any group $G$ the intersection of the kernels of morphisms from $G$ to groups in $\mathcal{K}$:
@@
More generally, consider a subgroup functor $\kappa$ for a class $\mathcal{C}$. We can then define a new subgroup functor $\theta_{(\mathcal{C},\kappa)}$ defined by
@@
It is easy to check that the obtained subgroup functor is indeed a radical.
For any group $G$ we will now abbreviate $G/\kappa(G)$ by $G/\kappa$.
For the following lemma we do not suppose that $\kappa(G/\kappa(G))=\{1\}$.
@@
Suppose that $\kappa$ is a radical. For any group we have that
@@
Let $\kappa$ be a radical. Let $(G,S)$ be a marked group. Let $R$ be a set of relations. Suppose that the relations of $R$ hold in $(G/\kappa,S)$. Then there is an isomorphism of marked groups:
@@ 4.2 First Lemma : Detecting stabilization on non-terminals is sufficient
Let $\kappa$ denote a radical. Let $n$ be a natural number. If
@@
Because $\kappa$ defines a radical, the composition $\pi\circ f$ factors through $\langle A|\bigcup_{i=0}^{n}\Phi^{i}(w_{0})\rangle/\kappa$. But by hypothesis
@@ 4.3 Second Lemma: Effective Noetherianity
In our applications, we will use the following radicals $\kappa$:
Nilpotent Quotients Metabelian quotients $\kappa(G)=\gamma_{k}(G)$, for a fixed $k$ in $\mathbb{N}$ $\kappa(G)=[[G,G],[G,G]]$ Finite quotients Residually free image $\kappa(G)=\bigcap_{\phi:G\rightarrow F}\ker(\phi),$ where $F$ is the given finite group $\kappa(G)=\bigcap_{\phi:G\rightarrow\mathbb{F}_{2}}\ker(\phi)$ Hyperbolic quotients $\kappa(G)=\bigcap_{\phi:G\rightarrow H}\ker(\phi),$ where $H$ is the given hyperbolic group
We include two effective noetherianity lemmas: one deals with nilpotent, metabelian and finite quotients, and the other one with free and hyperbolic quotients. The reason for this is that the stabilization that occurs is different in both cases: in the first three cases, when we quotient a free group by the considered characteristic subgroup, we get a group which satisfies the Max-n property: all increasing sequences of normal subgroups stabilize. And thus the sequence
@@
$\mathbb{F}_{A}/\langle\langle\Phi^{1}(w_{0})\rangle\rangle/\kappa\rightarrow\mathbb{F}_{A}/\langle\langle\Phi^{1}(w_{0})\rangle\rangle/\kappa/\langle\langle\Phi^{2}(w_{0})\rangle\rangle/\kappa\rightarrow...$
Lemma 58 (Effective noetherianity for Nilpotent, Metabelian and finite quotients).
Suppose that $\kappa$ is one of the subgroup functors associated to Nilpotent, Metabelian and finite quotients as presented in the table above. There exists an algorithm that on input a word $w_{0}$ and a set $\Phi$ of endomorphisms of a free group $\mathbb{F}_{A}$ produces a natural number $n$ such that
@@
This follows directly from the fact that in each case $\mathbb{F}_{A}/\kappa$ satisfies the Max-n property. This is well known for nilpotent groups. For metabelian groups, it is a theorem of Hall [Hal54]. For the subroup functor $\kappa$ associated to computing finite quotients, it is trivial, since in this case the group $\mathbb{F}_{A}/\kappa(\mathbb{F}_{A})$ is finite.
@@
This equality is equivalent to the fact that each $\phi$ in $\Phi$ already defines a group homomorphism in the quotient $(\mathbb{F}_{A}/\kappa(\mathbb{F}_{A}))/\langle\langle\bigcup_{i=0}^{n}\Phi^{i}(w_{0})\rangle\rangle.$ It is possible to check algorithmically that this is the case: it suffices to check that each of the finitely many elements of $\Phi^{n+1}(w_{0})$ is already the identity in $(\mathbb{F}_{A}/\kappa(\mathbb{F}_{A}))/\langle\langle\bigcup_{i=0}^{n}\Phi^{i}(w_{0})\rangle\rangle.$ This is decidable because the word problem is uniformly solvable for the three considered families of groups: nilpotent, metabelian, finite. ∎
We will now establish the corresponding result for free and hyperbolic quotients. A group $G$ is called equationally Noetherian if any infinite system of equations over $G$ is equivalent to a finite subsystem. What we will use to compute free quotients is the following theorem of Guba:
@@
Corollary 61.
If $H$ is a hyperbolic group, every residually-$H$ group is finitely presented as a residually-$H$ group.
Let $H$ be a hyperbolic group. Denote by $\kappa$ the subgroup functor associated to “residually-$H$ images”: $\kappa(G)=\bigcap_{\phi:G\rightarrow H}\ker(\phi)$ An equivalent formulation of the above corollary is:
Corollary 62.
@@
The abstract existence of $n$ such as in the lemma is given by Corollary 62. We thus have to justify that we can detect that for a certain $n$ the equality
@@
Lemma 64 (Effective noetherianity for computing hyperbolic quotients).
Let $H$ be a hyperbolic group, and $\kappa$ the associated radical. There exists an algorithm that on input a word $w_{0}$ and a set $\Phi$ of endomorphisms of a free group $\mathbb{F}_{A}$ produces a natural number $n$ such that
@@
The proof is identical to the case of free groups: we use the fact that $H$ is equationally noetherian [WR20], and that its universal theory is decidable [DG10] (and this uniformly in a given finite presentation of $H$). ∎

### 09 Apr 2024 v2

@@ 2.2.3 Tietze Transformations
Here, we give, thanks to Tietze transformations, natural conditions on a family of languages LAN so that the set of marked groups $\pi_{M}(\text{LAN})$ be saturated for the abstract group isomorphism relation. The conditions are given in Lemma 28.
@@
If LAN is a family of languages stable under taking homomorphic images and taking union with finite sets, then the set of marked groups that have a presentation in LAN is closed under group isomorphism: if a group has a presentation in LAN with respect to a marking, it has one with respect to any marking.
@@
Suppose that LAN is a family of uniformly recursively enumerable languages. Suppose furthermore that LAN is effectively stable under taking homomorphic images and taking union with finite sets. Then there is a procedure that takes as input a LAN-presentation of a marked groups $(G,S)$, and outputs a sequence of LAN-presentations of $G$ that contains at least one LAN-presentation for each marking of $G$.
@@
the hypotheses on the family LAN of languages immediately imply that this presentation is again a LAN-presentation. ∎
@@ 2.3 Computing quotients and dismissing infinite presentations
Let LAN be a family of languages, and $\mathcal{C}$ a set of finitely presented groups. In this section, we give conditions under which the following two problems are equivalent:
@@
Let $\mathcal{C}$ be a recursively enumerable class of finitely presented groups. Let LAN be a family of uniformly recursively enumerable languages. If the marked quotient problem in $\mathcal{C}$ is uniformly semi-decidable for groups given by presentations in LAN, then there is an algorithm which, given as input a presentation in LAN for a marked group in $\mathcal{C}$, produces a finite presentation for this group.
@@
An open problem is to find a family LAN of formal languages for which the above lemma can be applied with $\mathcal{C}$ being the set of all finitely presented groups. It is an open problem to decide whether groups given by EDT0L presentations have uniformly semi-decidable marked quotient problem for finitely presented groups.
@@
Let $\mathcal{C}$ be a class of finitely presented groups. Suppose that the isomorphism problem is solvable in $\mathcal{C}$ for groups given by finite presentations, and that $\mathcal{C}$ is recursively enumerable via finite presentations. Let LAN be family of languages. If the marked quotient problem in $\mathcal{C}$ is uniformly semi-decidable for groups given by presentations in LAN, then the isomorphism problem is solvable for groups in $\mathcal{C}$ given by presentations in LAN.
@@
Let $\mathcal{C}$ be a class of finitely presented groups stable under taking quotients. Let LAN be a family of languages for which the operation “taking union with finite sets” is computable. Then, if the marked isomorphism problem is solvable for groups in $\mathcal{C}$ with presentations in LAN, the $\mathcal{C}$-marked quotient problem is uniformly semi-decidable for groups given by LAN presentations.
@@ 3.1 Identification result
Proof.
From the definitions it is obvious that L-presentations produce their relators exactly by the DTF0L+FIN mechanism, with an additional constraint: the iterated morphisms should be group endomorphisms, and not simply monoid endomorphisms. Thus having an L-presentation implies having a DTF0L+FIN presentation. And of course, since DTF0L+FIN is a proper subfamily of EDT0L, having a presentation in DTF0L+FIN implies having a presentation in EDT0L.
@@ 3.2 Application to the marked isomorphism problem
We use the fact that equality of EDT0L languages is not semi-decidable, see [Roz72]. (Note that the theorem stated in [Roz72] is “equality is undecidable for EDT0L languages”. But for every family of uniformly recursive languages the equality problem is co-semi-decidable, since to prove that two languages are different it suffices to produce an element which belongs to one but not to the other. Hence the statement that equality is not decidable shows actually that equality is not semi-decidable for EDT0L languages.)
@@
One could expect much stronger results than the ones above. It would be very interesting to find examples of a family of languages LAN where the equality and inclusion problems are semi-decidable, while the isomorphism and quotient problem are not semi-decidable for groups with presentations in LAN.
@@
4 Unified approach to computing marked quotients
In this section we give a unified approach to computing marked quotients of EDT0L presented groups, thereby unifying and generalizing [BEH08] and [Har11]. (Note also that [BEH08] and [Har11] deal with L-presentations, and while this does not make a difference in terms of groups, it possibly does in terms of marked groups, and the following results deal with marked groups.) The following theorems were known:
@@ 4 Unified approach to computing marked quotients
Theorem 47 (Abelian-by-$k$-nilpotent quotients).
There is an algorithm that, given an EDT0L presentation for a marked group $(G,S)$, and an integer $k$, produces a finite presentation in the category of abelian-by-k-nilpotent groups of $(G/[\gamma_{k}(G),\gamma_{k}(G)],S)$.
What we mean in the above is that $(G/[\gamma_{k}(G),\gamma_{k}(G)],S)$ is given by a finite presentation together with the identity of abelian-by-k-nilpotent groups as an additional infinite set of relations.
@@
Theorem 49 (Quotients that are virtual direct products of hyperbolic groups).
There is an algorithm that, on input an EDT0L presentation for a marked group $(G,S)$, a finite presentation for a marked group $(H,S^{\prime})$ which is virtually a direct product of hyperbolic groups, and a function $f:S\rightarrow S^{\prime}$, decides whether or not $f$ can be extended to a group homomorphism.
The theorems above in each case rely on two properties :
  • For nilpotent and finite groups (the cases that were known), on noetherianity and uniform solvability of the word problem.
  • For abelian-by-nilpotent groups, noetherianity is replaced by the Max-n property [Hal54], which is noetherianity for normal subgroups. We also use uniform solvability of the word problem.
  • For direct products of hyperbolic groups, we rely on equational noetherianity instead of actual noetherianity, and on decidability of the universal theory instead of the word problem ([CHR20], extending the result of [DG10, Corollary 9.1] on hyperbolic groups). (Decidability of the universal Horn theory would suffice.)
@@ 4.1 Subgroup functors of fully characteristic subgroups
A residual is a subgroup functor $\kappa$ such that:
@@
Residuals are not unrelated the notion of “residually-$\mathcal{C}$” group, where $\mathcal{C}$ is a class of groups.
Lemma 53.
Suppose that $\kappa$ is a residual. Then there is a class $\mathcal{C}$ of groups such that for all $G$,
@@
And thus $G/\kappa(G)$ is the biggest residually-$\mathcal{C}$ quotient of $G$.
Proof.
We just take $\mathcal{C}$ to be $\{G,\kappa(G)=\{1\}\}$. Then:
$w\not\in\kappa(G)$ $\implies w\neq 1\text{ in $G/\kappa(G)$}$
@@
And:
$w\in\kappa(G)$ $\implies\forall H\in\mathcal{C},\forall\phi:G\rightarrow H,\phi(w)=1$
[…]
Note also the following corollary, which expresses the fact that the commutation property
$\forall G,\forall R\subseteq G,\,G/\kappa(G)/\langle\langle R\rangle\rangle=G/\langle\langle R\rangle\rangle/\kappa(G/\langle\langle R\rangle\rangle)$
characterizes group varieties.
Corollary 54.
Suppose that $\kappa$ is a residual, and that the property $\kappa(G)=\{1\}$ is quotient stable. Then the set of groups $G$ with $\kappa(G)=\{1\}$ is a group variety, and for any group $\kappa(G)$ is the subgroup of $G$ generated by the verbal subgroups associated to this variety. (It is thus itself a verbal subgroup when the variety is defined by finitely many identities.)
Proof.
This simply follows from the fact that a group variety is a set of groups stable by unrestricted direct product, taking subgroups and taking quotients, by a result of Birkhoff [Bir35]. Because the set of residually $\mathcal{C}$-groups, for any $\mathcal{C}$, is always stable by taking direct products and subgroups, quotient stability is the only missing hypothesis. ∎
For any group $G$ we will now abbreviate $G/\kappa(G)$ by $G/\kappa$. Also, for $R$ a set of elements of a group $G$, we denote by $G/R$ the quotient $G/\langle\langle R\rangle\rangle$, omitting to express the normal closure.
@@
Suppose that $\kappa$ is a residual. For any group we have that
@@
Let $\kappa$ be a residual. Let $(G,S)$ be a marked group. Let $R$ be a set of relations. Suppose that the relations of $R$ hold in $(G/\kappa,S)$. Then there is an isomorphism of marked groups:
@@ 4.2 First Lemma : Detecting stabilization on non-terminals is sufficient
Let $\kappa$ denote a residual. Let $n$ be a natural number. If
@@
Because $\kappa$ defines a residual, the composition $\pi\circ f$ factors through $\langle A|\bigcup_{i=0}^{n}\Phi^{i}(w_{0})\rangle/\kappa$. But by hypothesis
@@ 4.3 Second Lemma: Effective Noetherianity
In our applications, we will use the following residuals $\kappa$:
Nilpotent Quotients Abelian-by-Nilpotent Quotients $\kappa(G)=\gamma_{k}(G)$, for a given $k$ in $\mathbb{N}$ $\kappa(G)=[\gamma_{k}(G),\gamma_{k}(G)]$, for a given $k$ in $\mathbb{N}$ Residually free image Virtual direct products of hyperbolic groups $\kappa(G)=\bigcap_{\phi:G\rightarrow\mathbb{F}_{2}}\ker(\phi)$ $\kappa(G)=\bigcap_{\phi:G\rightarrow H}\ker(\phi),$ where $H$ is the given group
We include two effective noetherianity lemmas: one deals with abelian-by-nilpotent and finite quotients, and the other one with free and hyperbolic quotients. The reason for this is that the stabilization that occurs is different in both cases: in the first two cases, when we quotient a free group by the considered characteristic subgroup, we get a group which satisfies the Max-n property: all increasing sequences of normal subgroups stabilize. And thus the sequence
@@
$\mathbb{F}_{A}/\langle\langle w_{0}\rangle\rangle/\kappa\rightarrow\mathbb{F}_{A}/\langle\langle\Phi^{1}(w_{0})\rangle\rangle/\kappa/\langle\langle\Phi^{2}(w_{0})\rangle\rangle/\kappa\rightarrow...$
Lemma 59 (Effective noetherianity for abelian-by-nilpotent and finite quotients).
Suppose that $\kappa$ is one of the subgroup functors associated to abelian-by-nilpotent or finite quotients as presented in the table above. There exists an algorithm that on input a word $w_{0}$ and a set $\Phi$ of endomorphisms of a free group $\mathbb{F}_{A}$ produces a natural number $n$ such that
@@
This follows directly from the fact that in each case $\mathbb{F}_{A}/\kappa$ satisfies the Max-n property. It is a theorem of Hall [Hal54] for abelian-by-nilpotent groups. For the subroup functor $\kappa$ associated to computing finite quotients, it is trivial, since in this case the group $\mathbb{F}_{A}/\kappa(\mathbb{F}_{A})$ is finite.
@@
This equality is equivalent to the fact that each $\phi$ in $\Phi$ already defines a group homomorphism in the quotient $(\mathbb{F}_{A}/\kappa(\mathbb{F}_{A}))/\langle\langle\bigcup_{i=0}^{n}\Phi^{i}(w_{0})\rangle\rangle.$ It is possible to check algorithmically that this is the case: it suffices to check that each of the finitely many elements of $\Phi^{n+1}(w_{0})$ is already the identity in $(\mathbb{F}_{A}/\kappa(\mathbb{F}_{A}))/\langle\langle\bigcup_{i=0}^{n}\Phi^{i}(w_{0})\rangle\rangle.$ This is decidable because the word problem is uniformly solvable for the considered families of groups: abelian-by-nilpotent groups [BCM81] and finite groups. ∎
We will now establish the corresponding result for free quotients and quotients that are virtually direct products of hyperbolic groups. A group $G$ is called equationally Noetherian if any infinite system of equations over $G$ is equivalent to a finite subsystem. What we will use to compute free quotients is the following theorem of Guba:
@@
Note also that being equationally Noetherian is stable under finite direct products (this is immediate), and is also inherited by finite extensions by [BMR97]. Thus groups that are virtually direct products of hyperbolic groups are equationally Noetherian. We will use the following:
Proposition 62 (Houcine, [Hou07]).
If $H$ is an equationally Noetherian group, every residually-$H$ group is finitely presented as a residually-$H$ group.
Let $H$ be an equationally Noetherian. Denote by $\kappa$ the residual associated to “residually-$H$ images”: $\kappa(G)=\bigcap_{\phi:G\rightarrow H}\ker(\phi)$ An equivalent formulation of the above proposition is:
Proposition 63.
@@
The abstract existence of $n$ such as in the lemma is given by Proposition 63. We thus have to justify that we can detect that for a certain $n$ the equality
@@
Lemma 65 (Effective noetherianity for computing marked quotients that are virtual direct products of hyperbolic groups).
Let $H$ be a finitely generated group with a finite index subgroup which is a direct product of hyperbolic groups, and $\kappa$ the associated residual. There exists an algorithm that on input a word $w_{0}$ and a set $\Phi$ of endomorphisms of a free group $\mathbb{F}_{A}$ produces a natural number $n$ such that
@@
The proof is identical to the case of free groups: we use the fact that $H$ is equationally Noetherian (as hyperbolic groups are by [WR20], and as equational noetherianity is stable by direct product and finite extension [BMR97]), and that its universal theory is decidable [CHR20] (and this uniformly in a given finite presentation of $H$). ∎